1. Use Table A (Z-curve), sketch the curve and shade in the area that answers the question.

a) What is P ( Z > -1.75) = __________
b) What is the P ( -1.96 < z < - 1.0 ) = _________

2. Use the t-distribution, sketch the curve and shade in the area that answers the question. a) What is the P ( t > - 2.583 ) = ___________ when n = 18
a) What is the P ( t < _________ ) = .96 when n = 12
Answer #1

1)

Solution :

Using standard normal table,

(a)

P(Z > -1.75) = 1 - P(Z < -1.75) = 1 - 0.0401 = 0.9599

(b)

P(-1.96 < z < -1.0) = P(z < -1.0) - P(z < -1.96) = 0.1587 - 0.025 = 0.1337

2)

(a)

n = 18 degrees of freedom = n - 1 = 18 - 1 = 17

P(t > -2.583) = 0.9903

b)

n = 12

df = 12 - 1 = 11

P(t < ) = 0.96

P(t < 1.928) = 0.96